after selling a total of 427 adult and student tickets the cashier at a theater had receipts of $3263.50 if adult tickets sells for $8.50 and student tickets for 6.50 how many of each type of ticket were sold?

Answer by mananth(16946)

x Adult tickets
y Student stickets
Total tickets sold
x+y = 427.00 .............1
Value of tickets
8.50 x + 6.50 y = 3,263.50 .............2
Eliminate y
multiply (1)by -6.50
Multiply (2) by 1.00
-6.50 x -6.50 y = -2,775.50
8.50 x + 6.50 y = 3,263.50
Add the two equations
2.00 x = 488.00
/ 2.00
x = 244.00
plug value of x in (1)
x +y = 427.00
244.00 + y = 427.00
y = 427.00-244
y = 183.00
y = 183.00
244 Adult tickets
183 Student stickets
m.ananth@hotmail.ca